OpenStudy (anonymous):

The slope of the line passing through the points (2, 7) and (-4, 8) is -6 -1/2 -1/6

OpenStudy (anonymous):

m=y2-y1/x2-x1 Plug in the coordinates (2,7) = (x1,y1) (-4,8)=(x2,y2) sooo. m=8-7/-4-2 m=-1/6

OpenStudy (anonymous):